OpenStudy (anonymous):

Using the fundamental theorem of calculus, find the derivative of F(x) = definite intergral from x=10 to x=x (tan theta)dtheta steps please I've been trying to catch my mistake

OpenStudy (zarkon):

there is like 1 step ;)

OpenStudy (anonymous):

\[d/dt\int tan\theta d\theta=Log[cos[x]-Log[Cos[10]]]\]= d/dt Log[Cos[x]]

OpenStudy (anonymous):

Tan[x]

OpenStudy (zarkon):

\[\frac{d}{dx}\int\limits_{a}^{x}f(\theta)d\theta=f(x)\]

OpenStudy (zarkon):

you have to be careful about what values of x works here

OpenStudy (anonymous):

Cos[10] is negative which makes the answer imaginary, so..

OpenStudy (zarkon):

it doesn't also the answer is \(\tan(x)\)